# Model Validation Practice in Banking: A Structured Approach

### v1

@@ 4.3 Periodic Testing and Revalidation
Periodic testing and revalidation are essential components of on-going monitoring to ensure that models remain fit for purpose over time. This involves formal, scheduled reviews of the model’s performance and may be triggered either by routine cycles (e.g., quarterly, annually) or by specific events, such as significant economic changes or regulatory updates.
@@
  • Documentation and Regulatory Reporting: During periodic testing and revalidation, banks must ensure that the process is well-documented. This includes maintaining records of any changes to the model, rationale for those changes, and evidence that the model continues to meet regulatory and business standards. In the context of regulatory guidance, such as SR 11-7 in the U.S. or CRD IV in Europe, banks are required to periodically report the results of their model validation efforts to regulators.
@@ 4.4 Adaptive Maintenance and Model Refresh
As part of on-going monitoring, models may need to undergo adaptive maintenance, which includes updating model parameters or retraining the model with new data to reflect recent trends. This ensures that the model remains relevant in evolving environments.

### v2

@@ 4.3 Periodic Testing and Revalidation
Periodic testing and revalidation are essential components of ongoing monitoring to ensure that models remain fit for purpose over time. This involves formal, scheduled reviews of the model’s performance and may be triggered either by routine cycles (e.g., quarterly, annually) or by specific events, such as significant economic changes or regulatory updates.
@@
  • Documentation and Regulatory Reporting: During periodic testing and revalidation, banks must ensure that the process is well-documented. This includes maintaining records of any changes to the model, rationale for those changes, and evidence that the model continues to meet regulatory and business standards. In the context of regulatory guidance, such as SR 11-7, banks are required to periodically report the results of their model validation efforts to regulators.
@@ 4.4 Adaptive Maintenance and Model Refresh
As part of ongoing monitoring, models may need to undergo adaptive maintenance, which includes updating model parameters or retraining the model with new data to reflect recent trends. This ensures that the model remains relevant in evolving environments.